# Influence of momentum confinement in the phase-space dynamics of the Kramers-Henneberger atom

### v1

@@ IV.2 Phase-space behavior of coherent superpositions of eigenstates
Here, we compare the phase-space behavior in the full and KH time-averaged potential at relevant times of the dynamics given by the color points in Fig. 5b), which represent maxima, minima and mid-points of the autocorrelation functions computed using the time-averaged and full KH Hamiltonian. The times associated with the mid-point of the autocorrelation function for the time-averaged potential correspond to a probability density equally distributed across the two wells of the KH averaged potential. These instances are signposted with blue and red dots for the time-averaged and the full potential, although these behaviors occur at different times. The orange (black) dots indicate the times for which the wave packet is localized on the left (right) wells.
@@
In the left column of Fig. 6 we show snapshots of the Wigner quasiprobability function evaluated after solving the TDSE [Eq. (8)] using the equally weighted coherent superposition of the two KH eigenstates as the initial condition for the KH time-averaged potential. The specific times chosen for those snapshots match the dots in the green curve in Fig. 5. For the time-averaged potential, the Wigner quasiprobability distributions exhibit a cyclic behavior matching the oscillations of the autocorrelation function, without any significant spill outside the region of interest. The quasiprobability flow extends spatially to over 30 a.u. and oscillates between the wells of the dichotomous potential, spilling beyond the classical separatrix (see white curves in the figure). However, it is confined to a region of small momenta, within the green equienergy curves. Figs. 6a) and c) show quasiprobability located mainly in the central region between the two wells, with the Wigner plot in a) being the mirror image of that in c) regarding the x-axis. In contrast, for Figs. 6b) and d) the quasiprobability density is mostly located around the wells, and they are the mirror image of each other about the p-axis.
@@
The central column of Fig. 6 exhibits the Wigner functions at the same instants of time as in the left column, but computed using the full time-dependent potential. Overall, we observe a lack of agreement in the results from the full and time-averaged potential. Although there is a cyclic motion for the quasiprobability flow corresponding to the full dynamics, the times for which the Wigner functions are located on the left, right and central regions are mismatched. For instance, the positive quasiprobability density in Fig. 6b) [Fig. 6d)] is located around the left (right) well, while its counterpart in Fig. 6b’)[Fig. 6d’)] is located on the right (left). Similarly, the positive quasiprobablity density in Fig. 6a) [Fig. 6c)] is located in the phase-space region below (above) the saddle determined by the separatrix, while its counterpart in Fig. 6a’)[Fig. 6c’)] is located above (below) it. Furthermore, the behavior is not purely cyclic, with quasiprobability tails spilling towards higher momentum regions.
@@
Therefore, it is physically more meaningful to compare the Wigner quasiprobability flow after the onset of stabilization. These results are displayed in the right column of Fig. 6, in which we have chosen snapshots matching the colored points in the orange curve in Fig. 5 b). For these snapshots, the main dynamics of the system in the time-averaged case are reflected in the full expression. Interestingly, the comparison between the left and right columns demonstrates that despite the effect of the turn-on ramp in the beginning, the main features of the KH atom are retained in the Wigner function albeit at a later time interval. In both cases, we can see that the wave packet is essentially trapped in this spatial region, moving from one well to the other. The Wigner function exhibits a clockwise movement, with its period, which can be inferred from the autocorrelation function shown in Fig. 5 b), depending on the well separation. This is somewhat similar to the case of a diatomic molecule, with in our case playing the role of the internuclear distance. This cyclic motion is roughly in phase for both the time-averaged and full potentials. For instance, in Figs. 6a) and a”) [Figs. 6c) and c”)], there is a strong positive quasiprobability density in the central region below (above) the saddle, while, in Figs. 6b) and b”) [Figs. 6d) and d”)], they are peaked around the left [right] well. Still, the full results also exhibit ionization, observed as fringes extending the quasiprobability density towards higher momenta and distorting the cyclic motion. Nonetheless, they seem to be less pronounced than those observed before the onset of stabilization. This is very evident comparing Figs. 6d’) and d”), but can also be observed in subtler ways inspecting Figs. 6c’) and c”), or Figs. 6b’) and b”).
@@
In Fig. 7, we illustrate the onset of stabilization and its implication on these fringes more systematically. The snapshots in the figure were chosen such that the Wigner functions exhibit strong peaks around the right well of the KH potential, but considering the fully time-dependent dynamics. The times in Figs. 7a) and b) were chosen before one and two periods associated with the KH eigenstates, and Fig. 7c) was calculated for a time just before the field turn-off. For the three times, the field amplitude is kept constant, meaning they are after the field turn-on and before its tun-off. The fringes that form a tail associated with ionization become fainter with time as shown by comparing the three panels in Fig. 7. This suppression of ionization, along with wave-packet localization in one of the wells, indicates that we have transitioned to the stabilization regime.
@@
Next, we investigate to what extent the KH approximation (7), characterized by the zeroth order term of the Fourier expansion (5) of the full potential, holds, and how that manifests in the context of our Wigner functions. This will be done by initially using the ground state of the KH atom or the coherent superposition of the two KH eigenstates and allowing them to evolve under the full KH potential up to the point when the system enters the stabilization regime. This instance can be identified by observing the time at which the dichotomous structure emerges in the atomic probability density. Then, we take the evolved wave packet at this time as the initial condition for the time-averaged KH potential. Thus, the starting points of both the time-averaged and full KH potentials in the Wigner function calculation align more closely at the onset of stabilization.
@@
In Fig. 8 we show the results of this test for when we initially start off with a coherent superposition of KH eigenstates in the full KH potential expression and select the wave packet at a.u. as the initial condition for the time-averaged potential. The left and right columns of Fig. 8 have been computed for the time-averaged and full potential, respectively. The localization of the electron and the cyclic motion of the distribution over time are well described in the time-averaged case when comparing the same times. However, the tails associated with ionization disappear at later times if the time propagation is performed using the time-averaged potential. This can be seen by comparing Figs. 8b) and b’), and Figs. 8c) and c’), and means that ionization takes place even if the stabilization regime has been reached.
@@ V Conclusions
We started by understanding the Wigner quasiprobability distributions associated with the KH eigenstates, and comparing them with the classically determined equienergy curves. For a time-averaged KH potential, these Wigner functions are stationary and have well-defined parities. The ground-state Wigner function is localized mostly at the origin, between the two wells, and is non-negative, while the excited-state Wigner function is peaked around the wells. There is also a large negativity near the origin, which hints at the nonclassical behavior of the Wigner functions in that region. A common feature is that, to some extent, they are within the constraints posed by the equienergy curves of the time-averaged KH potential, roughly following their shapes. Furthermore, an inspection of the classical equienergy curves shows that there is no spatial confinement beyond the separatrix, but that the momentum ranges are restricted and very small. This relates to the phase-space studies using classical methods [21, 40, 41], and the computation of bounds for the quantum mechanical ionization probability [42, 43], which refer to vanishing drift momentum as a pre-requisite for stabilization.
@@
A wave packet constructed using an equally weighted coherent superposition of these eigenstates will exhibit a cyclic motion, which resembles that observed for molecules in our previous publications [29, 49]. The frequency associated with the motion is proportional to the energy difference between the two KH eigenstates, and is captured by the periodic oscillations in the corresponding autocorrelation function. One should note, however, that, for molecules in intense fields, a cyclic motion associated with the Wigner quasiprobability densities only occurs for internuclear separations much smaller than the distance between the wells of the KH potential. This is caused by a confinement effect, which, in phase space, is related to nested separatrices. For intermediate internuclear distances, these separatrices open, and enhanced ionization occurs until, eventually, for internuclear distances comparable to those in the present work, the different wells in the molecule will behave independently. This means that the quasiprobability flow between them will be strongly suppressed.
@@
Therefore, the analogy between the KH atom and a molecule is limited in scope. For molecules, there is tunnel ionization from one center to the other and spatial confinement, while, for the KH atom, the energy ranges associated with the wave packet and with the KH eigenstates are above that of the separatrix, suggesting over-the-barrier motion. There is a lack of spatial confinement by a potential well, but the nearly vanishing drift momentum leads to a localized and cyclic wave packet in the stabilization regime. In the phase space, this is related to the equienergy curves not closing for the KH potential, but, instead, confining the motion to the region of small momenta. Thus, our phase-space analysis elucidates that confinement for the KH atom occurs in momentum space rather than position space. Recently, the lack of spatial confinement in the KH atom being a limitation for its analogy with a molecule has been highlighted in the context of pump-probe schemes [15].
@@
Finally, an interesting question would be whether the level of suppression of ionization and spatial localization that we have observed is good enough for building a theoretical model of a qubit. The feasibility of this approach will depend on the robustness of this localization, precise laser field control and how well the electron’s motion can be confined and controlled in a scalable way without leading to excessive decoherence. This will form the basis of future investigations.

### v2

@@ IV.2 Phase-space behavior of coherent superpositions of eigenstates
Here, we compare the phase-space behavior in the full and KH time-averaged potential at relevant times of the dynamics given by the color points in Fig. 6(b), which represent maxima, minima and mid-points of the autocorrelation functions computed using the time-averaged and full KH Hamiltonian. The times associated with the mid-point of the autocorrelation function for the time-averaged potential correspond to a probability density equally distributed across the two wells of the KH averaged potential. These instances are signposted with blue and red dots for the time-averaged and the full potential, although these behaviors occur at different times. The orange (black) dots indicate the times for which the wave packet is localized on the left (right) wells. The corresponding Wigner quasiprobability distributions are plotted in Fig. 7, for the time-averaged KH dynamics [left column] and full temporal evolution [middle and right columns].
@@
In the left column of Fig. 7 we show snapshots of the Wigner quasiprobability function evaluated after solving the TDSE [Eq. (8)] using the equally weighted coherent superposition of the two KH eigenstates as the initial condition for the KH time-averaged potential. The specific times chosen for those snapshots match the dots in the green curve in Fig. 6(b). For the time-averaged potential, the Wigner quasiprobability distributions exhibit a cyclic behavior matching the oscillations of the autocorrelation function, without any significant spill outside the region of interest. The quasiprobability flow extends spatially to over 30 a.u. and oscillates between the wells of the dichotomous potential, spilling beyond the classical separatrix (see white curves in the figure). However, it is bounded to a region of small momenta, within the green equienergy curves. Figs. 7(a) and (c) show quasiprobability located mainly in the central region between the two wells, with the Wigner plot in (a) being the mirror image of that in (c) regarding the x-axis. In contrast, for Figs. 7(b) and (d) the quasiprobability density is mostly located around the wells, and they are the mirror image of each other about the p-axis.
@@
The central column of Fig. 7 exhibits the Wigner functions at the same instants of time as in the left column, that is, indicated by the colored points in the green curve in Fig. 6(b), but computed using the full time-dependent potential. Overall, we observe a lack of agreement in the results from the full and time-averaged potential. Although there is a cyclic motion for the quasiprobability flow corresponding to the full dynamics, the times for which the Wigner functions are located on the left, right and central regions are mismatched. For instance, the positive quasiprobability density in Fig. 7(b) [Fig. 7(d)] is located around the left (right) well, while its counterpart in Fig. 7(b’)[Fig. 7(d’)] is located on the right (left). Similarly, the positive quasiprobablity density in Fig. 7(a) [Fig. 7(c)] is located in the phase-space region below (above) the saddle determined by the separatrix, while its counterpart in Fig. 7(a’)[Fig. 7(c’)] is located above (below) it. Furthermore, the behavior is not purely cyclic, with quasiprobability tails spilling towards higher momentum regions.
@@
Therefore, it is physically more meaningful to compare the Wigner quasiprobability flow after the onset of stabilization. These results are displayed in the right column of Fig. 7, in which we have chosen snapshots matching the colored points in the orange curve in Fig. 6(b). For these snapshots, the main dynamics of the system in the time-averaged case are reflected in the full expression. Interestingly, the comparison between the left and right columns demonstrates that despite the effect of the turn-on ramp in the beginning, the main features of the KH atom are retained in the Wigner function albeit at a later time interval. In both cases, we can see that the wave packet is essentially trapped in this spatial region, moving from one well to the other. The Wigner function exhibits a clockwise movement, with its period, which can be inferred from the autocorrelation function shown in Fig. 6(b), depending on the well separation. This is somewhat similar to the case of a diatomic molecule, with in our case playing the role of the internuclear distance. This cyclic motion is roughly in phase for both the time-averaged and full potentials. For instance, in Figs. 7(a) and (a”) [Figs. 7(c) and (c”)], there is a strong positive quasiprobability density in the central region below (above) the saddle, while, in Figs. 7(b) and (b”) [Figs. 7(d) and (d”)], they are peaked around the left [right] well. Still, the full results also exhibit ionization, observed as fringes extending the quasiprobability density towards higher momenta and distorting the cyclic motion. Nonetheless, they seem to be less pronounced than those observed before the onset of stabilization. This is very evident comparing Figs. 7(d’) and (d”), but can also be observed in subtler ways inspecting Figs. 7(c’) and (c”), or Figs. 7(b’) and (b”).
@@
In Fig. 8, we illustrate the onset of stabilization and its implication on these fringes more systematically, reinforcing the statement that the tail-shaped features are associated with ionization. The snapshots in the figure were chosen such that the Wigner functions exhibit strong peaks around the right well of the KH potential, but considering the fully time-dependent dynamics. The times in Figs. 8(a) and (b) were chosen before one and two periods associated with the KH eigenstates, and Fig. 8(c) was calculated for a time just before the field turn-off. For the three times, the field amplitude is kept constant, meaning they are after the field turn-on and before its turn-off. The fringes that form a tail towards higher momentum regions become fainter with time as shown by comparing the three panels in Fig. 8. This suppression of ionization, along with wave-packet localization in one of the wells, indicates that we have transitioned to the stabilization regime.
@@
Next, we investigate to what extent the KH approximation (7), characterized by the zeroth order term of the Fourier expansion (5) of the full potential, holds, and how that manifests in the context of our Wigner functions. This will be done by initially using the ground state of the KH atom or the coherent superposition of the two KH eigenstates and allowing them to evolve under the full KH potential up to the point when the system enters the stabilization regime. This instance can be identified by observing the time at which the dichotomous structure emerges in the atomic probability density [see Fig. 3 and the ensuing discussion]. Then, we take the evolved wave packet at this time as the initial condition for the time-averaged KH potential. Thus, the starting points of both the time-averaged and full KH potentials in the Wigner function calculation align more closely at the onset of stabilization.
@@
In Fig. 9 we show the results of this test for when we initially start off with a coherent superposition of KH eigenstates in the full KH potential expression and select the wave packet at a.u. as the initial condition for the time-averaged potential. The left and right columns of Fig. 9 have been computed for the time-averaged and full potential, respectively. The localization of the electron and the cyclic motion of the distribution over time are well described in the time-averaged case when comparing the same times. However, the tails associated with ionization disappear at later times if the time propagation is performed using the time-averaged potential. This can be seen by comparing Figs. 9(b) and (b’), and Figs. 9(c) and (c’), and means that ionization takes place even if the stabilization regime has been reached.
@@ V Conclusions
We started by understanding the Wigner quasiprobability distributions associated with the KH eigenstates, and comparing them with the classically determined equienergy curves. For a time-averaged KH potential, these Wigner functions are stationary and have well-defined parities. The ground-state Wigner function is localized mostly at the origin, between the two wells, and is non-negative, while the excited-state Wigner function is peaked around the wells. There is also a large negativity near the origin, which hints at the nonclassical behavior of the Wigner functions in that region. A common feature is that, to some extent, they are within the constraints posed by the equienergy curves of the time-averaged KH potential, roughly following their shapes. Furthermore, an inspection of the classical equienergy curves shows that there is no spatial confinement beyond the separatrix, but that the momentum ranges are restricted and very small. This is specific to the atomic potential used in this work, as we have verified that for a soft-core potential with the parameters used in [42], there will be a larger momentum spread for Wigner functions coherent superpositions of KH eigenstates (not shown). Still, the momentum regions occupied by the quasiprobability flow is bounded, with vanishing net drift momentum. This relates to the phase-space studies using classical methods [21, 42, 43], and the computation of bounds for the quantum mechanical ionization probability [44, 45], which refer to vanishing drift momentum as a pre-requisite for stabilization.
@@
A wave packet constructed using an equally weighted coherent superposition of these eigenstates, will exhibit a cyclic motion, which resembles that observed for molecules in our previous publications [30, 48]. The frequency associated with the motion is proportional to the energy difference between the two KH eigenstates, and is captured by the periodic oscillations in the corresponding autocorrelation function. Although the initial coherent superposition of the KH eigenstates is similar to what is obtained by combining molecular eigenstates of different parity, and also leads to similar effects (cyclic motion and population trapping) to that of a molecule, the physical mechanisms behind both are different. One should note, that for molecules in intense fields, a cyclic motion associated with the Wigner quasiprobability densities only occurs for internuclear separations much smaller than the distance between the wells of the KH potential. This is caused by a confinement effect, which, in phase space, is related to nested separatrices. For intermediate internuclear distances, these separatrices open, and enhanced ionization occurs until, eventually, for internuclear distances comparable to those in the present work, the different wells in the molecule will behave independently. This means that the quasiprobability flow between them will be strongly suppressed.
@@
Therefore, the analogy between the KH atom and a molecule is limited in scope. For molecules, there is tunnel ionization from one center to the other and spatial confinement, while, for the KH atom, the energy ranges associated with the wave packet and with the KH eigenstates are above that of the separatrix, suggesting over-the-barrier motion. There is a lack of spatial confinement by a potential well, but the momentum remains bounded, so that a cyclic motion is observed. Specifically for the potential studied here, the nearly vanishing drift momentum leads to a localized and cyclic wave packet in the stabilization regime. In the phase space, this is related to the equienergy curves not closing for the KH potential, but, instead, representing a boundary for the motion, which remains within the region of small momenta. Other atomic potentials may lead to larger instantaneous momenta and a more convoluted motion for the quasiprobability flow, but the momentum remains bounded if stabilization has been attained. Recently, the lack of spatial confinement in the KH atom being a limitation for its analogy with a molecule has been highlighted in the context of pump-probe schemes [15].
@@
Finally, an interesting question would be whether the level of suppression of ionization and spatial localization that we have observed is good enough for building a theoretical model of a qubit. The feasibility of this approach will depend on the robustness of this localization, precise laser field control and how well the electron’s motion can be kept in bounded phase-space regions and controlled in a scalable way without leading to excessive decoherence. This will form the basis of future investigations.